OpenStudy (anonymous):

K is the midpoint of JL , JL= 4x-2 and Jk= 7. Find x , KJ , and JL

OpenStudy (anonymous):

4x-2=7 4x=7+2 4x=9 9/4=2.5 x=2.5 4(2.5)-2=7 JL=7